# Undergraduate Course Allocation

### v1

@@ 2 Pseudo-Markets with Priorities
    $x_{s}^{*}=\max_{\left(\succsim_{s}\right)}\left\{x^{\prime}_{s}\subseteq{\cal C}:\sum_{c\in{\cal C}}p_{c,r_{s,c}}^{*}x_{c}^{\prime}\leq b_{s}^{*}\right\}$
@@
    where $\sum_{\{s\in{\cal S}:r_{s,c}=r^{*}_{c}\}}x^{*}_{s,c}>0$ if $\sum_{s\in{\cal S}}x^{*}_{s,c}>0$ and $\overline{b}=1+\beta+\epsilon$ for some small $\epsilon>0$.
@@
    The market-clearing error is less than $\alpha$: $||z^{*}||_{2}\leq\alpha,$ where $z^{*}=(z_{1}^{*},....,z_{M}^{*})$ and:
@@
Unlike Budish (2011), the above definition of an approximate competitive equilibrium allows for course prices to depend on priority levels, making the vector of prices $p^{*}=\left\{p^{*}_{c,r}\right\}_{c\in{\cal C},r\in\{1,...,R\}}\in\mathbb{R}^{M\!R}$. Condition ($\ast$2) ensures that each course has a cutoff level of priority, where higher priority students obtain seats in the course for free, students at the cutoff level of priority face non-negative prices, and lower priority students cannot afford a seat. In doing so, condition ($\ast$2) guarantees no students are allocated a seat in a course when a higher priority student who would like a seat does not receive one. A similar condition appeared in He et al. (2018) in the context of random allocation with priorities.
Requiring that $\sum_{\{s\in{\cal S}:r_{s,c}=r^{*}_{c}\}}x^{*}_{s,c}>0$ if $\sum_{s\in{\cal S}}x^{*}_{s,c}>0$ ensures course $c$ has some students at the cutoff level demand seats, if there is any demand. If we allow $r^{*}_{c}$ to satisfy $\sum_{\{s\in{\cal S}:r_{s,c}=r^{*}_{c}\}}x^{*}_{s,c}=0$, raising the cutoff level of priority to $r^{*}_{c}+1$ does not change any student’s demand for the course. In turn, choosing $r^{*}_{c}$ so that $\sum_{\{s\in{\cal S}:r_{s,c}=r^{*}_{c}\}}x^{*}_{s,c}>0$ guarantees that the cutoff level of priority cannot be raised without removing a student from the course.
@@
The market-clearing error for a course depends only on its price at $r=1$, the lowest priority level. Under-enrollment is only counted as an error if the price at level $r=1$ is positive. This requirement is a non-trivial extension of Budish (2011)’s definition to settings with course priorities. Alternatively, if a course’s market-clearing error depends on the price at the cutoff priority level, one can change the cutoff level and artificially lower the error. For example, consider an under-enrolled course with a price of zero at and above the cutoff priority level and a price of $\overline{b}$ below the cutoff. By this definition, the market-clearing error is zero. However, the unfilled course seats should be counted towards the market-clearing error, as they may be eliminated if the price below the cutoff level is decreased.
@@
Budish (2011) establishes the existence of an approximate competitive equilibrium from equal incomes for any budget inequality $\beta>0$ with an upper bound on the market-clearing error of $\alpha=\sqrt{kM/2}$. [^8] Budish (2011) shows the existence of an approximate equilibrium from equal incomes with a market-clearing error at most $\sqrt{\min\{2k,M\}M}/2$. As we assume $k\leq M/2$, the upper bound reduces to $\sqrt{kM/2}$. This bound is proportional to the square root of the dimension of the price space. In the presence of $R$ priority levels for each of the $M$ courses, the dimension of the price space is $M\!R$. This could potentially require a large market-clearing error for a Pseudo-Market Equilibrium with Priorities to exist. The main idea of the proof of Theorem 1 is to reduce the effective dimension of the price space from $M\!R$ to $M$, where we can establish the existence of a Pseudo-Market Equilibrium with Priorities. In particular, for any vector $t\in\mathbb{R}^{M}$, we can define priority-specific prices $p\in\mathbb{R}^{M\!R}$ by setting:
@@
for each course $c\in{\cal C}$ and priority level $r\in\{1,...,R\}$. This parameterization guarantees that, for any $t\in[0,R\overline{b})$, the corresponding price vector $p$ satisfies condition ($\ast$2). We then prove the existence of an approximate equilibrium in the $M$-dimensional space.
@@
Theorem 1 establishes the existence of a Pseudo-Market Equilibrium with Priorities, extending the main result of Budish (2011) to environments with course priorities. This equilibrium has the same upper bound on the market-clearing error as the approximate competitive equilibrium from equal incomes in the same environment without priorities. The proof of Theorem 1 is presented in Appendix A.
@@ 2.1 The Pseudo-Market with Priorities Mechanism
    Each student $s$ is assigned a random budget $b^{*}_{s}$ drawn uniformly from $[1,1+\beta]$.
@@
    Compute an $(\alpha,\beta)$-Pseudo-Market Equilibrium with Priorities $(x^{*},p^{*},b^{*})$. Allocate courses to students according to $x^{*}$.
In general, the PMP mechanism depends on the level of allowable market-clearing error $\alpha$ and the level of budget inequality $\beta$. For results that hold for all non-negative $\alpha\geq 0$ and $\beta\geq 0$, we avoid this dependence in our exposition.
@@
Every Pseudo-Market with Priorities mechanism results in an allocation that is approximately stable.
The proof is postponed to Appendix A. The main reason that the outcome of the PMP mechanism is not stable is the possibility of over- and under-enrolled courses. Stability requires a feasible allocation and only allows a course to have unfilled seats if no student can benefit from taking one. The PMP mechanism can violate these conditions when the market-clearing error is nonzero. However, as Theorem 1 shows, such instances are rare.
@@
Consider a setting with two students ${\cal S}=\{1,2\}$ and two courses ${\cal C}=\{A,B\}$, with $q_{A}=q_{B}=k=1$. Student preferences are $\{A\}\succ_{1}\{B\}\succ_{1}\varnothing$ and $\{B\}\succ_{2}\{A\}\succ_{2}\varnothing$. Student budgets are $b^{*}_{1}=1$ and $b^{*}_{2}=1+\beta$ for some $10<\beta<$ 1. Course priorities are the opposite of student preferences, with $r_{1,B}=r_{2,A}=2$ and $r_{1,A}=r_{2,B}=1$. Budgets $b^{*}$, cutoff priority levels $r^{*}_{A}=r^{*}_{B}=2$, priority-specific prices $p^{*}_{A}=p^{*}_{B}=(2,1)$, and allocation:
@@
Every Pseudo-Market with Priorities mechanism results in an allocation that is approximately priority-constrained efficient.
@@
Theorem 2 prevents the possibility that a student can benefit from taking a seat in a course that a student with lower priority obtains. However, priorities in undergraduate course allocation are weak, with hundreds of students often at the same level of priority across many courses. The next result addresses the question of fairness among students whose priorities are weakly ordered in terms of their priorities across all courses.
@@
The result of Theorem 4 extends the one established in Budish (2011) for settings where all students are on the same priority level. With priorities based on seniority and student department, the above result implies that the PMP mechanism with a small budget inequality produces an allocation in which any two students from the same department and the same year of study have schedule envy bounded by a single course toward one another. [^12] This result can be strengthened. The PMP mechanism results in an allocation $x^{*}$ such that for any two students $s,s^{\prime}\in\cal S$ such that $r_{s,c}\geq r_{s^{\prime},c}$ for each $c\in x^{*}_{s^{\prime}}$, envy is bounded by a single course. In particular, rather than all courses, $s$ needs only to be at a weakly higher priority level in the courses being taken by $s^{\prime}$.
@@
Last, we relate the PMP mechanism to some special cases discussed in previous literature. In the case of no priorities, our mechanism reduces to the approximate competitive equilibrium from equal incomes mechanism of Budish (2011). On the other hand, courses may have strict priorities over students. In this case, we have a two-sided matching problem. As students can view sets of courses as complements, no stable matching is guaranteed to exist. Still, Theorem 1 confirms a Pseudo-Market Equilibrium with Priorities exists with a small error. So, our mechanism does not necessarily coincide with any stable matching mechanism.
@@
Finally, He et al. (2018) show how to extend the pseudo-market approach of Hylland and Zeckhauser (1979) to school choice settings with random allocations. Their pricing structure ensures that their pseudo-market equilibrium allocation is fair. We adopt a similar pricing structure but analyze deterministic allocations rather than random allocations. While deterministic allocations are preferable from a practical market design point of view, they are associated with additional complications in the existence of an equilibrium.
@@ 3.1 Student and Course Data
Notes: We use capital letters to denote college names.
We utilize university data containing the final allocation of courses to students from the Spring 2018 semester at a private institution in the mid-Atlantic region. There are 6 02360236\,0236 023 students divided into four years of study and $756$ courses. The university is made up of seven colleges. Colleges are the largest constituent units of the university and typically contain several departments. For example, the Department of Economics and the Department of English are housed in the College of Liberal Arts at many U.S. universities. [^15] We only consider undergraduate students. We merge fifth-year and fourth-year students, as there are few fifth-year students in the data, and both groups are treated equally in the course registration process. Some graduate and exchange students take undergraduate courses, but we exclude these students. Colleges differ in the size of their student population and the number and type of courses they offer, as shown in Table 1. Some colleges serve the whole university, offering general education courses in math and science, whereas others mainly offer courses only for their students.
Course enrollments. Table 2 shows that $11.2\%$ of all courses are at or above their listed capacity. This percentage serves as an important indicator of overlap in student interest and the need for a proper mechanism to allocate seats in over-demanded courses. Moreover, the table indicates that the university allows a significant number of courses to be over-enrolled (around $7.3\%$). The listed maximum capacity for such courses is intended to ensure proper class dynamics and balanced course sections. For example, teaching business communication or drama courses of a large size is difficult, but these courses can often accommodate a few students above their listed capacity. This fits our treatment of capacity constraints in the Pseudo-Market with Priorities mechanism, which allows enrollment above a course’s capacity. However, in some courses, the maximum capacity is strictly binding. These situations are typically associated with the physical limitations of the classrooms. We set each course’s capacity equal to the maximum of the capacity observed in the data and the true enrollment.
Course seat reservations. The data records course seat reservations, each specifying a course and a number of seats reserved for students in certain departments and years of study. The concept of course seat reservations is somewhat different from the concept of a priority structure used in the theoretical part of the paper. Rather than saving a fixed number of seats for specific groups of students, a priority structure ensures that certain student groups have a claim on course seats over other student groups. Table 3 presents quantiles of the distributions of course capacities, enrollment, and total number of reserved seats from the university data. In total, 33 4553345533\,45533 455 course seats were available across courses, 23 3692336923\,36923 369 seats were occupied by students, and 13 9221392213\,92213 922 were reserved at the start of the course registration process.
@@
Notes: The table reports information about over-enrollment of courses in the university data. Maximum capacity is a soft constraint on enrollment for many courses.
@@
Course priorities. Using the data on course seat reservations, we create a course priority structure that reflects the university administration’s intentions. [^16] There is a tradeoff between using a course priority structure and using course seat reservations. Reservations allow the university administration to specify distributional concerns by reserving a certain amount of a course’s seats to multiple, different groups of students without prioritizing one of the other. However, they require specifying an exact number of reserved seats, which is inflexible in environments with uncertain course demand. With priorities, if there are few high priority students, the seats are free to be allocated to students of lower priority. However, priorities do not address distributional concerns. See Celebi and Flynn (2024) for an innovative analysis merging priorities and reserves into the concept of adaptive priorities. In practice, the course registration process closely follows a Random Serial Dictatorship with course reserves, where students select courses in order of seniority (with ties broken randomly) and courses make seat reservations based on department and year of study. Hence, the process favors students in later years of study by giving them earlier time slots to choose courses, resembling year-specific priorities (see Section 1). In addition, courses reserve seats for students in specific departments and years of study, resembling department-specific priorities. [^17] Universities that follow this process include Princeton, Johns Hopkins, Duke, Vanderbilt, Washington University in St. Louis, Columbia, Notre Dame, and Carnegie Mellon. However, each university operates slightly differently. Within a year of study, many universities (e.g., Duke) randomly stagger students across time slots rather than having all register at once. In addition, some universities (e.g., Vanderbilt) reserve seats for students in specific majors or degree programs rather than departments or years of study. We combine these notions by using a hybrid priority structure in our simulations, with year-specific priorities taking precedence over department-specific priorities. We analyze the case in which department-specific priorities take precedence over year-specific priorities in Appendix D. [^18] There are more than two ways how the year-specific and department-specific priorities can be combined.
Example 2 provides an illustration for Real Analysis in Table 4. We drop the exact number of reserved seats and assign the highest priority to 4th-year students in Department 2, followed by 4th-year students in other departments, 3rd-year students in Department 2, and 3rd-year students in other departments. No seats are reserved for 1st- and 2nd-year students, making all 2nd-year and all 1st-year students the bottom two levels.
@@
Student utilities. The university data does not contain information on students’ preferences over course schedules. To recover this information, we use a parametric form for student utilities and the method of simulated moments to calibrate its parameters (see, e.g., Agarwal, 2015). We assume that each student can take up to $k=5$ courses and that students’ preferences are additive over courses. [^19] Some students are enrolled in six or even seven courses in the data. These students account for less than $3\%$ of the total student population. Student $s$’s utility from taking course $c$ is denoted by $u_{sc}$ and has the following form:
$u_{sc}={\color[rgb]{0,0,0}Y^{\prime}_{s}\theta}+Y^{\prime}_{s}\Gamma H_{c}+{\color[rgb]{0,0,0}z_{c}}+\varepsilon_{sc},$ (1)
where $Y_{s}$ and $H_{c}$ are binary vectors of observed student and course characteristics, respectively, $\theta$ is a vector of unobserved coefficients, $\Gamma$ is a matrix of unobserved coefficients conformable with vectors $Y_{s}$ and $H_{c}$, $z_{c}$ is an unobserved course-specific component, and $\varepsilon_{sc}$ is random utility component that is assumed to have a normal distribution $\varepsilon_{sc}\sim N(0,\sigma^{2})$ with unobserved parameter $\sigma$. As normalizations, the value of the outside option from not taking a course is $u_{s0}=0$ and the standard deviation of the noise parameter is $\sigma=1$.
@@
The parametric form in equation (1) contains three non-random components. The first, $Y^{\prime}_{s}\theta$, is a horizontal component determined by the student’s characteristics. The sole nonzero element in $Y_{s}$ indicates student $s$’s college $a\in{\cal A}=\{A,...,G\}$ and student $s$’s year of study $y\in{\cal Y}=\{1,...,4\}$. That is, $Y_{s}\in\{0,1\}^{|{\cal A}||{\cal Y}|}$ and $Y_{s}^{\prime}\theta=\theta_{ay}$, where vector $\theta\in\mathbb{R}^{|{\cal A}||{\cal Y}|}$ has $|{\cal A}||{\cal Y}|=28$ elements to calibrate. The larger this horizontal component is, the more courses student $s$ wants to enroll in. This accounts for variability in the number of courses taken by students from different colleges and years of study (see Table C.1 in Appendix C).
@@
The second component, $Y^{\prime}_{s}\Gamma H_{c}$, is an interaction term that determines the heterogeneity of students’ preferences over courses. We restrict the matrix $\Gamma$ to have $49$ independent parameters $\gamma_{aa^{\prime}}$, where $a$ is student $s$’s college and $a^{\prime}$ is course $c$’s college. This accounts for large heterogeneity in patterns of course enrollment across colleges (see Table C.2 in Appendix C). The third component, ${\color[rgb]{0,0,0}z_{c}}$, is a vertical component that measures the popularity or quality of course $c$. The vector $z$ has length $756$, with one entry for each course.
@@
We cannot jointly identify the parameters $\theta$, $\Gamma$, and $z$, as they enter in equation (1) in an additive way. We normalize the average horizontal component across students in the same college to $\overline{\theta}_{a}=0$ and set the interaction terms $\gamma_{aa}$ to 00 for each $a\in{\cal A}$. These normalizations allow for an interpretation of $z_{c}$ as the average student utility from taking course $c$ among students in the same college as course $c$. The value of the parameter $\gamma_{aa^{\prime}},a\neq a^{\prime}$ is the average increase in student utility from taking course $c^{\prime}$ in college $a^{\prime}$ over taking course $c$ in student’s own college $a$ when both courses have the same values of vertical components (i.e., $z_{c}=z_{c^{\prime}}$). We provide more details on the parameter normalization on p. C.1 in the Appendix C. Overall, we have $819$ independent parameters $\theta$, $\gamma$, and $z$ to calibrate.
@@
Last, we assume that each student’s choice set is limited to $80$ courses. This assumption is motivated by the observation that any given student typically only considers taking a subset of all offered courses in a given term and is made for computational purposes (see also Budish and Cantillon, 2012; Diebold and Bichler, 2017). We randomly draw these $80$ courses among those taken by at least one student from the same college and year. The probability that a student in college $a$ and year $y$ draws course $c$ is proportional to the share of seats taken in course $c$ by students in college $a$ and year $y$, among all seats taken by such students. [^20] We expand some students’ choice sets to make sure that courses have enough reserve-eligible students with the course in their choice sets. This ensures that our calibration process runs smoothly. We elaborate on this on p. C.1 in Appendix C. Not all $80$ utility entries are necessarily positive.
@@
We use the method of simulated moments to identify the model’s parameters. A description of the calibration procedure, interpretations of the calibrated parameter values, and model fit results are left to Appendix C. Here, we discuss two major complications that arose during the calibration process. The first obstacle in recovering student utilities was that, in practice, course reservations do not remain constant during the registration process. After the first week of course registration, student demand for courses becomes more or less clear. To ensure complete course enrollment, departments start relaxing course reservations by admitting students from waiting lists. While there are best practices for handling waiting lists, this phase of enrollment is done solely at each department’s discretion. [^21] The best practices include prioritizing more senior students and students in the course’s department. As a result, a course’s final enrollment can violate its seat reservations set at the beginning of the process. We use Hall’s Marriage Theorem to identify course reserve violations and minimally adjust the number of reserved seats. The seat reservations are lowered in 170 of the 756 courses to ensure the number of reserved seats is consistent with the allocation observed in the data.
The second obstacle concerned calibrating the vertical components $z_{c}$ for courses at maximum capacity. For such courses, no value of $z_{c}$ can lead to a higher enrollment than we observe in the data. While we could have used modern econometrics techniques to partially identify course fixed effects and provide inference on parameters using moment inequalities (see Chernozhukov, Chetverikov and Kato, 2019), it is not clear which parameter values from the partially identified set to choose for the student utility model. Instead, we identified the vertical components for courses at maximum capacity using data on how quickly the seats in those courses were occupied during the course enrollment process. Appendices C.2-C.3 provide a detailed discussion of both obstacles.

### v2

@@ 2 Pseudo-Markets with Priorities
    $x_{s}^{*}=\max_{\left(\succsim_{s}\right)}\left\{x^{\prime}_{s}\subseteq{\cal C}:\sum_{c\in{\cal C}}p_{c,r_{s,c}}^{*}x_{s,c}^{\prime}\leq b_{s}^{*}\right\}$
@@
    where $\overline{b}=1+\beta+\epsilon$ for some small $\epsilon>0$.
@@
    The market-clearing error is less than $\alpha$: $z^{*}||_{2}\leq\alpha,$ where $z^{*}=(z_{1}^{*},....,z_{M}^{*})$ and:
@@
Unlike Budish (2011), the above definition of an approximate competitive equilibrium allows for course prices to depend on priority levels, making the vector of prices $p^{*}=\left\{p^{*}_{c,r}\right\}_{c\in{\cal C},r\in\{1,...,R\}}\in\mathbb{R}^{M\!R}$. Condition ($\ast$2) ensures that each course has a cutoff level of priority, where higher priority students obtain seats in the course for free, students at the cutoff level of priority face some price, and lower priority students cannot afford a seat. In doing so, condition ($\ast$2) guarantees no students are allocated a seat in a course when a higher priority student who would like a seat does not receive one. A similar condition appeared in He et al. (2018) in the context of random allocation with priorities.
@@
The market-clearing error for a course depends only on whether its price equals to zero at $r=1$, the lowest priority level. Under-enrollment is only counted as an error if the price at level $r=1$ is positive. This requirement is a non-trivial extension of Budish (2011)’s definition to settings with course priorities. Alternatively, if a course’s market-clearing error depends on whether the price at the cutoff priority level is zero, one can change the cutoff level and artificially lower the error. For example, consider an under-enrolled course with a price of zero at and above the cutoff priority level and a price of $\overline{b}$ below the cutoff. By the alternative definition, the market-clearing error is zero. However, the unfilled course seats should be counted towards the market-clearing error, as they may be eliminated if the price below the cutoff level is decreased.
@@
Budish (2011) establishes the existence of an approximate competitive equilibrium from equal incomes for any budget inequality $\beta>0$ with an upper bound on the market-clearing error of $\alpha=\sqrt{kM/2}$. [^8] Budish (2011) shows the existence of an approximate equilibrium from equal incomes with a market-clearing error at most $\sqrt{\min\{2k,M\}M}/2$. As we assume $k\leq M/2$, the upper bound reduces to $\sqrt{kM/2}$. This bound is proportional to the square root of the dimension of the price space. In the presence of $R$ priority levels for each of the $M$ courses, the dimension of the price space is MRM\!R. This could potentially require a large market-clearing error for a Pseudo-Market Equilibrium with Priorities to exist. The main idea of the proof of Theorem 1 is to reduce the effective dimension of the price space from MRM\!R to $M$, where we can establish the existence of a Pseudo-Market Equilibrium with Priorities. In particular, for any vector $t\in\mathbb{R}^{M}$, we can define priority-specific prices $p\in\mathbb{R}^{M\!R}$ by setting:
@@
for each course $c\in{\cal C}$ and priority level $r\in\{1,...,R\}$. This parameterization guarantees that, for any $t\in[0,R\overline{b})^{M}$, the corresponding price vector $p$ satisfies condition ($\ast$2). We then prove the existence of an approximate equilibrium in the $M$-dimensional space.
@@
Theorem 1 establishes the existence of a Pseudo-Market Equilibrium with Priorities with the same upper bound on the market-clearing error as the approximate competitive equilibrium from equal incomes, extending the main result of Budish (2011) to environments with course priorities. The proof of Theorem 1 is presented in Appendix A.
@@ 2.1 The Pseudo-Market with Priorities Mechanism
    Each student $s$ is assigned a random budget $b^{*}_{s}$ drawn uniformly from $1,1+\beta]$.
@@
    Compute an $(\alpha,\beta)$-Pseudo-Market Equilibrium with Priorities $(x^{*},p^{*},b^{*})$ using an anonymous method. Allocate courses to students according to $x^{*}$.
In general, the PMP mechanism depends on the level of allowable market-clearing error $\alpha$ and the level of budget inequality $\beta$. We avoid this dependence in our exposition for results that hold for all non-negative $\alpha\geq 0$ and $\beta\geq 0$.
@@
The Pseudo-Market with Priorities mechanism results in an allocation that is approximately stable.
The proof is postponed to Appendix A. The main reason that the outcome of the PMP mechanism is not stable is the possibility of over- and under-enrolled courses. Stability requires a feasible allocation and only allows a course to have unfilled seats if no student can benefit from taking one. The PMP mechanism can violate these conditions when the market-clearing error is nonzero. However, Theorem 1 suggests instances will be rare.
@@
Consider a setting with two students ${\cal S}=\{1,2\}$ and two courses ${\cal C}=\{A,B\}$, with $q_{A}=q_{B}=k=1$. Student preferences are $\{A\}\succ_{1}\{B\}\succ_{1}\varnothing$ and $\{B\}\succ_{2}\{A\}\succ_{2}\varnothing$. Student budgets are $b^{*}_{1}=1$ and $b^{*}_{2}=1+\beta$ for some $0<\beta<1$. Course priorities are the opposite of student preferences, with $r_{1,B}=r_{2,A}=2$ and $r_{1,A}=r_{2,B}=1$. Budgets $b^{*}$, cutoff priority levels $r^{*}_{A}=r^{*}_{B}=2$, priority-specific prices $p^{*}_{A}=p^{*}_{B}=(2,1)$, and allocation:
@@
The Pseudo-Market with Priorities mechanism results in an allocation that is approximately priority-constrained efficient.
@@
Theorem 2 prevents the possibility that a student can benefit from taking a seat in a course that a student with lower priority obtains. However, priorities in undergraduate course allocation are weak, with hundreds of students often at the same level of priority across many courses. The next result addresses fairness between students whose priorities are weakly ordered in the same way across all courses.
@@
Theorem 4 extends the result established in Budish (2011) for settings without priorities. With priorities based on seniority and student department, the above result implies that, with a sufficiently small budget inequality, the PMP mechanism produces an allocation in which any two students from the same department and the same year of study have schedule envy bounded by a single course toward one another. [^12] This result can be strengthened. The PMP mechanism results in an allocation $x^{*}$ such that for any two students $s,s^{\prime}\in\cal S$ such that $r_{s,c}\geq r_{s^{\prime},c}$ for each $c\in x^{*}_{s^{\prime}}$, envy is bounded by a single course. In particular, rather than all courses, $s$ needs only to be at a weakly higher priority level in the courses being taken by $s^{\prime}$.
@@
We conclude this section by considering special cases discussed in previous literature. In the case of no priorities, our mechanism reduces to the approximate competitive equilibrium from equal incomes mechanism of Budish (2011). On the other hand, courses may have strict priorities over students. In this case, we have a two-sided matching problem. As students can view sets of courses as complements, no stable matching is guaranteed to exist. Still, Theorem 1 confirms a Pseudo-Market Equilibrium with Priorities exists with a small error. So, our mechanism does not necessarily coincide with any stable matching mechanism.
@@
Allowing random allocations, He et al. (2018) show how to extend the pseudo-market approach of Hylland and Zeckhauser (1979) to school choice settings. Their pricing structure ensures that their pseudo-market equilibrium allocation is fair. We adopt a similar pricing structure but analyze deterministic allocations rather than random allocations. While deterministic allocations are preferable from a practical market design point of view, they are associated with additional complications in equilibrium existence.
@@ 3.1 Student and Course Data
Notes: We use capital letters to denote college names. There are $1565$, $1611$, $1422$, and $1425$ students in the first, second, third, and fourth year of study, respectively.
We utilize data from the Spring 2018 semester at a private institution in the mid-Atlantic region, covering 6 0236\,023 students across four years of study and $756$ courses across seven colleges (see Table 1). Colleges are the largest constituent units of the university and typically contain several departments. For example, the Department of Economics is housed in the College of Liberal Arts at a typical U.S. university. Colleges differ in the size of their student population and the number and type of courses they offer. Some colleges serve the whole university, whereas others mainly offer courses only for their students. [^15] We exclude graduate and exchange students. Fourth- and fifth-year students are treated as one group.
@@
Notes: Maximum capacity is a soft constraint on enrollment for many courses.
Course enrollments. The university allows courses to be over-enrolled. Table 2 shows that $11.2\%$ of courses meet or exceed their listed capacity, with $7.3\%$ of courses exceeding it. The listed maximum capacity in over-enrolled courses is intended to ensure proper class dynamics and balanced course sections. This fits our treatment of capacity constraints in the Pseudo-Market with Priorities mechanism, which allows over-enrollment through the market-clearing error. We set each course’s capacity equal to the maximum of the capacity observed in the data and the true enrollment.
Course seat reservations. In practice, the course registration process follows a Random Serial Dictatorship with course reserves, where students select courses in order of seniority (with ties broken randomly) and courses make seat reservations for students in certain departments and years of study. Seat reservations allow courses to express priorities for students, ensuring that students in specific groups have access to a fixed number of a course’s seats. [^16] See Celebi and Flynn (2024) for a framework analyzing both reservations and priorities. Table 3 presents quantiles of the distributions of course capacities, enrollment, and total number of reserved seats from the university data. 33 45533\,455 course seats were available across courses, 23 36923\,369 seats were occupied, and 13 92213\,922 were reserved at the start of course registration.
@@
Using the data on seat reservations, we create a priority structure that reflects the university’s intentions. The registration process favors more senior students by giving them earlier time slots to choose courses, resembling year-specific priorities. In addition, courses reserve seats for students in specific departments and years of study, resembling department-specific priorities. [^17] Universities that follow these priorities include Princeton, Johns Hopkins, Duke, Vanderbilt, Washington University in St. Louis, Columbia, Notre Dame, and Carnegie Mellon. However, each university operates slightly differently. Within a year of study, many universities (e.g., Duke) randomly stagger students across time slots rather than having all register at once. In addition, some universities (e.g., Vanderbilt) reserve seats for students in specific majors or degree programs rather than departments or years of study. We combine these notions into a hybrid priority structure, with year-specific priorities taking precedence over department-specific priorities. Example 2 provides an illustration for Real Analysis in Table 4. As no seats are reserved for 1st- and 2nd-year students, there are no students at the fourth and second levels of priority.
@@
Student utilities. The data does not contain information on student preferences over course schedules. To recover this information, we use a parametric form for student utilities and the method of simulated moments to calibrate its parameters (see, e.g., Agarwal, 2015). We assume that each student can take up to $k=5$ courses and that preferences are additive over courses. [^18] Less than 3% of the student body is enrolled in more than five courses in the data. Student $s$’s utility from taking course $c$, $u_{sc}$, takes the form
$u_{sc}={\color[rgb]{0,0,0}Y^{\prime}_{s}\theta}+Y^{\prime}_{s}\Gamma H_{c}+{\color[rgb]{0,0,0}z_{c}}+\varepsilon_{sc}.$ (1)
@@
The parametric form in equation (1) contains three non-random components. The first, $Y^{\prime}_{s}\theta$, is a horizontal component determined by the student’s characteristics. $Y_{s}$ is a binary vector that indicates student $s$’s college $a\in{\cal A}=\{A,...,G\}$ and year of study $y\in{\cal Y}=\{1,...,4\}$. So, $Y_{s}^{\prime}\theta=\theta_{ay}$, where vector $\theta\in\mathbb{R}^{|{\cal A}||{\cal Y}|}$ has $|{\cal A}||{\cal Y}|=28$ elements to calibrate. The larger $\theta_{ay}$ is, the more courses student $s$ wants to enroll in. This accounts for variability in the number of courses taken by students from different colleges and years of study (see Table C.1 in Appendix C).
@@
The second component, $Y^{\prime}_{s}\Gamma H_{c}$, is an interaction term that determines the heterogeneity of students’ preferences over courses. $H_{c}$ is a binary vector of observed course characteristics indicating course $c$’s college. We restrict matrix $\Gamma$ to have only $49$ independent parameters $\gamma_{aa^{\prime}}$, where $a$ is student $s$’s college and $a^{\prime}$ is course $c$’s college. This accounts for patterns of course enrollment across colleges (see Table C.2 in Appendix C).
The third component, $\color[rgb]{0,0,0}z_{c}$}, is a vertical component that measures the popularity or quality of course $c$. The vector $z$ has length $756$, with one entry for each course. The last term $\varepsilon_{sc}$ is a random utility component that is assumed to have a normal distribution $\varepsilon_{sc}\sim N(0,\sigma^{2})$ with unobserved parameter $\sigma$. As normalizations, the value of the outside option from not taking a course is $u_{s0}=0$ and the standard deviation of the noise parameter is $\sigma=1$.
@@
We cannot jointly identify the parameters $\theta$, $\Gamma$, and $z$ without additional normalization as they enter in equation (1) in an additive way. Hence, we normalize the average horizontal component across students in the same college to $\overline{\theta}_{a}=0$ and set the interaction terms $\gamma_{aa}$ to 0 for each $a\in{\cal A}$. These normalizations allow for an interpretation of $z_{c}$ as the average student utility from taking course $c$ among students in the same college as course $c$. The value of the parameter $\gamma_{aa^{\prime}},a\neq a^{\prime}$ is the average increase in student utility from taking course $c^{\prime}$ in college $a^{\prime}$ over taking course $c$ in student’s own college $a$ when both courses have the same values of vertical components (i.e., $z_{c}=z_{c^{\prime}}$). Overall, we have $819$ independent parameters $\theta$, $\gamma$, and $z$ to calibrate.
@@
Last, we assume that each student’s choice set is limited to $80$ courses. This assumption is motivated by the observation that any given student typically only considers taking a subset of all offered courses in a given term and is made for computational purposes. We randomly draw these $80$ courses among those taken by at least one student from the same college and year. The probability that a student in college $a$ and year $y$ draws course $c$ is proportional to the share of seats taken in course $c$ by students in college $a$ and year $y$, among all seats taken by such students. [^19] We expand some students’ choice sets to make sure enough reserve-eligible students are interested in each course. This ensures that our calibration process runs smoothly. We elaborate on this in Appendix C. Not all $80$ utility entries are necessarily positive.
@@
We use the method of simulated moments to identify the model’s parameters. A description of the calibration procedure, interpretations of the calibrated parameter values, and model fit results are left to Appendix C. Here, we discuss two obstacles in the calibration process. First, in practice, course reservations do not remain constant during the registration process. After the first week of registration, demand for courses becomes more or less clear. To ensure complete course enrollment, departments start relaxing course reservations by admitting students from waiting lists. While there are best practices for handling waiting lists, this phase of enrollment is done solely at each department’s discretion. As a result, a course’s final enrollment can violate its seat reservations set at the beginning of the process. We use Hall’s Marriage Theorem to identify course reserve violations and minimally adjust the number of reserved seats. The seat reservations are lowered in 170 of the 756 courses to ensure the number of reserved seats is consistent with the allocation observed in the data.
The second obstacle concerned calibrating the vertical components $z_{c}$ for courses at maximum capacity. For such courses, no value of $z_{c}$ can lead to a higher enrollment than observed in the data. We identified the vertical components for courses at maximum capacity using data on how quickly seats in those courses were occupied during the registration process.